# “It’s Always a Losing Game”: How Workers Understand and Resist Surveillance Technologies on the Job

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 2025 ACM Conference on Computer-Supported Cooperative Work and Social Computing (CSCW); October 18–22, 2024; Bergen, Norwaybooktitle: CSCW’25: ACM Conference on Computer-Supported Cooperative Work and Social Computing, October 18–22, 2025, Bergen, Norwayprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
What do warehouse workers, delivery workers, and call center workers have in common? Their productivity is tracked and compared to the average of other workers. When workers are slower than average, they face disciplinary action. This causes workers to sacrifice work quality and safety to improve their times. When everyone does this, they compete against more unrealistic expectations. It’s always a losing game. Anyone can make a tracking app, but in the end it’s poor management because it doesn’t take into account worker burnout and safety. This harms both worker and employer.” — Post on Reddit [R141]111”Throughout this paper, Reddit posts have been paraphrased or slightly altered to protect worker anonymity
As the above Reddit post articulates, the now ubiquity of electronic monitoring, or Workplace Surveillance Technologies (WSTs), has transformed how workers are managed across industries. Often associated with blue-collar jobs in manufacturing and call centers since the 1990s, WSTs underwent a significant expansion in use during the COVID-19 pandemic (Kantor and Sundaram, 2022). This rise has seen these technologies flourish in previously unmonitored sectors, including higher-paying creative and technical professions. This shift has also ushered in an era of granular, real-time scrutiny for workers. A combination of ”old” technologies like cameras and ”new” technologies like artificial intelligence (AI) is now being deployed to track employee productivity, location, physical environment, biometrics, computer activity, and even personal data (Hickok and Maslej, 2023).
Paralleling this rise, there has been an increasing interest within CSCW and related scholarly communities in how data are collected about workers’ performance, as well as how these evaluations affect pay structures and one’s ability to sustain employment (Jarrahi et al., 2021; Lee et al., 2015). Although initially deployed within the context of gig work, scholars have charted how algorithmic management techniques are now prevalent across traditional workplace settings and employment relationships (Kellogg et al., 2020). Work has also examined how these data-driven management strategies are combined with other forms of monitoring (e.g., customer ratings) to generate an ever more detailed assessment of workers’ activities (Cameron, 2022; Rosenblat, 2016). Scholars such as Sannon et al. have documented how workers engage in risk mitigation strategies associated with such oversight, including self-protective surveillance behaviors (Sannon et al., 2022), while others describe the potential for collective sensemaking as a strategy for worker contestation (Holten Møller et al., 2021).
@@
Workers across our study describe how WSTs contribute to a culture of distrust in the workplace that harms both workers and managers alike. Beyond privacy concerns, workers identify these technologies as causing significant stress, reducing their productivity, and increasing their risk of disciplinary action. At the same time, they see these technologies as fostering paranoia and distrust among managers toward their employees. In response, some workers employ various resistance tactics, such as commiseration, obfuscation, soldiering, and quitting to help protect them from intrusive monitoring. However, the reliance on these more individualized tactics point to the limited power and protections workers have when subjected to highly-surveilled and fissured workplaces.
Our paper makes three contributions to the CSCW community. First, we offer an empirical study on how workers have been impacted by WSTs since COVID-19 and the rise of remote work across different sectors. Second, we broaden understanding of worker resistance by highlighting the myriad of tactics workers employ to navigate and circumvent surveillance and its relationship to building worker power. Third, we identify opportunities for CSCW researchers to support workers in these countermeasures and to help bridge the gap toward more collective action.
In what follows, we first describe the contemporary landscape of WSTs and their impacts on workers. We then discuss current understandings of worker resistance and how they build power in response to overt surveillance and poor working conditions, and the ways past HCI research has supported these efforts. Next, we detail our methods that guide our content analysis and interview process. Drawing on our findings that explore how workers experience and resist surveillance technologies, we demonstrate the need for HCI research to go beyond a focus on mitigating the risks of surveillance and support efforts in resistance and refusal to its use.
@@ 2.1. Workplace Surveillance Technologies
Technological advancements since the late twentieth century have led to the age of digital surveillance, making it easier for employers to constantly monitor their workers both inside and outside the workplace (Introna, 2000; Vitak and Zimmer, 2023a; Ajunwa et al., 2017). WSTs have the ability to track and even infer a worker’s productivity, behavior, and personal information through digital activity monitoring, camera and audio monitoring, and localization and biophysical monitoring (Ajunwa et al., 2017; Ball, 2010; Cousineau et al., 2023). As AI and sensing technologies become more affordable and accessible, workplaces have introduced more intrusive forms of surveillance that often involve an assemblage of networked devices to generate continuous fine-grained data about their workers (Marx, 2002). The responsibility of workplace monitoring has also become more distributed in nature, where employees are expected to aid in their own surveillance through participation in workplace wellness programs and productivity apps marketed towards improving their work and wellbeing (Ajunwa et al., 2017; Ball, 2010; Chowdhary et al., 2023; Kawakami et al., 2023; Adler et al., 2022; Saha and Iqbal, 2023; Ajunwa et al., 2016). The increased gigification of work, coupled with the rise of remote work during the COVID-19 pandemic, has seen a proliferation of these technologies being used beyond the workplace, further blurring the lines between work and private life (Calacci, 2022; Allyn, 2020; Satariano, 2020; Zickuhr, 2021; Gould et al., 2023).
@@
Although these proposed benefits highlight the potential for monitoring technologies to support workers, there have been numerous concerns that undermine this vision, including threats to worker privacy, autonomy, health, and safety, and the reinforcement of existing biases (Awumey et al., 2024; Constantinides and Quercia, 2022). These technologies often collect extraneous and irrelevant information that do not meaningfully correlate to attributes of success on the job (Roemmich et al., 2023; Pritchard et al., 2015). In particular, AI systems that capture and infer worker emotion, mood, and affect have been susceptible to inaccuracies and bias (Roemmich et al., 2023; Boyd and Andalibi, 2023; Corvite et al., 2023; Das Swain et al., 2023; Kaur et al., 2022). Workers have reported experiencing notable shifts in behavior and interpersonal dynamics, including increased stress, disciplinary action, distrust, and competition (Pritchard et al., 2015; Levy, 2015; Alimahomed-Wilson and Reese, 2021). Pritchard et al. (Pritchard et al., 2015) found that London bus drivers performed a form of “self-surveillance” that involved proactively altering their behavior to avoid being flagged by monitoring systems or to beat their previous performance scores. Power differentials and a lack of transparency make it difficult for workers to meaningfully consent to these technologies, as they have little power to opt out or be in positions to understand and control the extent of monitoring, their data, or how they are being assessed (Chowdhary et al., 2023; Sannon et al., 2022; Ajunwa, 2020; Ajunwa et al., 2016). In some cases, monitoring systems have led to unsafe working conditions and poorer performance when they force workers to perform extra labor to continuously check, negotiate, and respond to these systems (Pritchard et al., 2015; Lu et al., 2021; Sannon et al., 2022). Women and minority workers, in particular, have been disproportionately impacted as they are more likely to take on low-wage jobs and face specific biases that make them more susceptible to invasive monitoring (Ball, 2010; Stark et al., 2020; Anjali Anwar et al., 2021; Negró n and Nguyen, 2023).
@@
Workers have limited legal protections in the face of workplace surveillance (Zickuhr, 2021). In the US, there are currently no federal laws in place and legislation is limited in piecemeal form at the local and state levels (Ajunwa et al., 2017). The California Privacy Rights Act of 2020 is one of the few examples that offers some privacy protections to workers (Gray, 2020). Although frequently used as an exemplar for omnibus legislation that provides robust data protections, there remain gaps within the European Union’s GDPR that enables organizations to surveil workers while still being in compliance (Jodka, 2018; Horan, 2018). As a result, there have been calls for stronger organizational and governmental regulation to protect worker autonomy, privacy, and collective rights (Das Swain et al., 2023; Ebert et al., 2021; Nguyen, 2021; Calacci and Stein, 2023) and for technologists to take a worker-centered design approach that meaningfully engages with workers as collaborators, limits unnecessary data capture, and recognizes the nuanced and contextual aspects of their labor that are harder to measure or make visible (Calacci, 2022; Holten Møller et al., 2021; Zhang et al., 2022; Kawakami et al., 2023; Cameron et al., 2021; Akridge et al., 2024). Our work expands upon this literature by taking a more holistic view on how workers direct experience and make sense of surveillance technologies across sectors since the rise of remote work as a result of the COVID-19 pandemic by drawing from worker testimonials on Reddit and in-depth interviews.
@@ 2.2. Worker Power and Resistance
Since the beginning of workplace surveillance, there has been worker resistance (Sale, 1995; Luff, 2008). Earlier studies of WSTs have shown how call center workers manipulate their metrics, such as leaving lines open after hanging up or pretending to talk on the phone (Frenkel et al., 1998; Ball, 2010; Callaghan and Thompson, 2002). Despite technologies becoming more sophisticated, workers have still found creative ways to circumvent them. For example, gig workers use sousveillance (Mann and Ferenbok, 2013; Mann et al., 2002; Cecchinato et al., 2021) techniques to monitor “up” by tracking payment discrepancies in ridehail work or collecting information about specific requesters on Amazon Mechanical Turk (Griffith, 2022; Hara et al., 2018; Do et al., 2024; Sannon et al., 2022). Additionally, workers use self-surveillance techniques by tracking their own personal data as a way to hold their employers and customers accountable (Sannon et al., 2022; Gallagher et al., 2023). “Worker inquiry” through data collection and analysis has long been an advocacy tool used by workers and unions since the days of Taylorism (Sannon et al., 2022; Khovanskaya et al., 2019; United Electrical and of America, 1943; Khovanskaya and Sengers, 2019; Dalal, 2024; Alquati, 2019). Sannon et al. (Sannon et al., 2022) broadly call these tactics self-protective surveillance that help to protect workers from risky work situations. They outlined other resistance tactics gig workers used to protect their work and privacy, including using platforms like Reddit to commiserate with other workers, creating false personas, or using technologies like VPNs to protect their personal information (Sannon et al., 2022). Surveillance studies scholar Gary Marx outlines 11 resistance tactics in total that workers use to subvert surveillance, including discovery, avoidance, piggybacking, switching, distorting, blocking, masking, breaking, refusal, explaining and contesting, and cooperation (Marx, 2003). Other workers have collectively organized to form unions or participate in strikes to fight against overt surveillance and poor working conditions (Alimahomed-Wilson and Reese, 2021).
@@
To address these power asymmetries, some HCI researchers have investigated how technological tools can better support workers in these countermeasures (Hara et al., 2018; Salehi et al., 2015; Irani and Silberman, 2013; Calacci and Pentland, 2022; Zhang et al., 2023; Toxtli and Savage, 2023; Toxtli et al., 2021; Platform, [n. d.]; Cooperative, [n. d.]; Job, [n. d.]; Lee et al., 2015). One of the most well-known examples is Turkopticon, an online platform where Amazon Mechanical Turk workers can rate, review, and vet requestors. Calacci & Pentland (Calacci and Pentland, 2022) co-built the Shipt calculator with worker groups to allow gig workers to track and share wage data and surface pay disparities. While these two tools have helped support workers in collective data stewardship, sensemaking, and organizing, Calacci states that more can be done to support “Digital Workism” by co-building tools with workers that support their goals in collective organizing while preserving their autonomy, privacy, and wellbeing (Calacci, 2022, 2023). While past literature mainly focuses on the resistance tactics employed by gig workers, our paper examines the countermeasures used by workers across multiple types of employment, including remote workers, call center workers, and data workers.
@@ 3.1.1. Reddit as a Site of Study
In line with other studies that focus on sensitive topics such as mental health, addiction, and pregnancy loss, we chose to study forums on Reddit due to its support for anonymous disclosure (Ammari et al., 2018; Andalibi et al., 2016; Andalibi and Forte, 2018; van der Nagel and Frith, 2015; Gauthier et al., 2022), which helps people talk about sensitive or stigmatized topics without having to disclose their real identity (Newman et al., 2011; van der Nagel and Frith, 2015). Researchers also have looked to Reddit to understand the experiences of gig workers, since online communities have been crucial for gig workers to share knowledge and build community despite the distributed nature of their work (Sannon et al., 2022; Carlos Alvarez de la Vega et al., 2021; Ramesh et al., 2023). In particular, Sannon et al. used Reddit to understand gig workers’ experiences regarding privacy and surveillance (Sannon et al., 2022). Expanding upon this work, we chose Reddit as a site to understand workers’ experiences with surveillance technologies across different sectors and how they used the platform to share knowledge and support one another.
@@ 3.1.2. Data Collection
We first identified posts that discussed workplace surveillance through manual keyword searches using surveillance-related terms such as “track,” “monitor,” and “surveil,” paired with keywords like “work” and “job.” We then randomly selected 66 posts to do a preliminary round of coding to find other relevant keywords, initial codes, and identify worker-centered subreddits to include in our study. We excluded subreddits that did not primarily focus on workers, (e.g. r/Privacy, r/Technology, r/legaladvice), subreddits that focus on managers or professional roles that do surveillance-type work (e.g. r/sysadmin, r/askmanagers), as well as industry or company-specific subreddits (e.g r/doordash_drivers, r/mturk, r/Nanny). This allowed us to understand how information was being shared cross-sectorally between workers in different industries and to get a sense of the breadth of the different types of bossware and its commonalities and differences. Finally, we excluded 1 worker subreddit that had explicit policies against data collection. In total, we selected 9 subreddits, including 2 general work subreddits (r/jobs, r/work), 5 specific to remote work (r/overemployed, r/RemoteJobs, r/remotework, workfromhome, r/WorkOnline) and 2 that had a political lean (r/antiwork, r/WorkersStrikeBack). Based on our initial coding, we finalized our keyword list to the following: “track,” “monitor,” “surveil,” “spy,” “productivity,” “micromanage,” and “bossware.”
Since we were interested in understanding the impacts of COVID-19 on workplace surveillance, we only included posts written between 1/1/2019 and 12/31/2022. We also excluded posts that had less than average engagement, which was calculated by identifying the median number of comments per post according to each subreddit. Based on this criteria, we created our filtered dataset through Pushshift API222stuck_in_the_matrix, Watchful1, and RaiderBDev. Reddit comments/submissions 2005-06 to 2023-12. Academic Torrents. https://academictorrents.com/details/9c263fc85366c1ef8f5bb9da0203f4c8c8db75f4 , which contained a total of 9,362 filtered posts. For each post, we extracted the keywords used, their comments, date, title, username, URL, and text.

### v2

@@ Abstract.
copyright: acmlicensedjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 2article: CSCW004publicationmonth: 4doi: 10.1145/3710902ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
What do warehouse workers, delivery workers, and call center workers have in common? Their productivity is tracked and compared to the average of other workers. When workers are slower than average, they face disciplinary action. This causes workers to sacrifice work quality and safety to improve their times. When everyone does this, they compete against more unrealistic expectations. It’s always a losing game. Anyone can make a tracking app, but in the end it’s poor management because it doesn’t take into account worker burnout and safety. This harms both worker and employer.” — Post on Reddit [R141] [^1] Throughout this paper, Reddit posts have been paraphrased or slightly altered to protect anonymity.
As the above Reddit post articulates, the current ubiquity of electronic monitoring, or Workplace Surveillance Technologies (WSTs), has transformed how workers are managed across industries. Often associated with blue-collar jobs in manufacturing and call centers since the 1990s, WSTs underwent a significant expansion in use during the COVID-19 pandemic (Kantor and Sundaram, 2022). This rise has seen these technologies flourish in previously unmonitored sectors, including higher-paying creative and technical professions. This shift has also ushered in an era of granular, real-time scrutiny for workers. A combination of “old” technologies like cameras and “new” technologies like artificial intelligence (AI) is now being deployed to track employee productivity, location, physical environment, biometrics, computer activity, and even personal data (Hickok and Maslej, 2023).
Paralleling this rise, there has been an increasing interest within CSCW and related scholarly communities in how data are collected about workers’ performance, as well as how these evaluations affect pay structures and one’s ability to sustain employment (Jarrahi et al., 2021; Lee et al., 2015). Although initially deployed within the context of gig work, scholars have charted how algorithmic management techniques are now prevalent across traditional workplace settings and employment relationships (Kellogg et al., 2020). Research has also examined how these data-driven management strategies are combined with other forms of monitoring (e.g., customer ratings) to generate an ever more detailed assessment of workers’ activities (Cameron, 2022; Rosenblat, 2016). Scholars such as Sannon et al. (Sannon et al., 2022) have documented how workers engage in risk mitigation strategies associated with such oversight, including self-protective surveillance behaviors, while others describe the potential for collective sensemaking as a strategy for worker contestation (Holten Møller et al., 2021).
@@
Workers across our study describe how WSTs contribute to a culture of distrust in the workplace that harms both workers and managers alike. Beyond privacy concerns, workers identify these technologies as causing significant stress, reducing their productivity, and increasing their risk of disciplinary action. At the same time, they see these technologies as fostering paranoia and distrust among managers toward their employees. In response, some workers employ various resistance tactics, such as commiseration, obfuscation, soldiering, and quitting to help protect them from intrusive monitoring. However, the reliance on these more individualized tactics points to the limited power and protections workers have when subjected to highly-surveilled and fissured workplaces (Weil, 2014).
Our paper makes three contributions to the CSCW community. First, we offer an empirical study on how workers have been impacted by WSTs since COVID-19 and the rise of remote work across different sectors. Second, we broaden understanding of worker resistance by highlighting the myriad tactics workers employ to navigate and circumvent surveillance and its relationship to building worker power. Third, we identify opportunities for CSCW researchers to support workers in these countermeasures and to help bridge the gap toward more collective action.
In what follows, we first describe the contemporary landscape of WSTs and their impacts on workers. We then discuss current understandings of worker resistance and how they build power in response to overt surveillance and poor working conditions, and the ways past HCI research has supported these efforts. Next, we detail the methods that guide our content analysis and interview process. Drawing on our findings that explore how workers experience and resist surveillance technologies, we demonstrate the need for HCI research to go beyond a focus on mitigating the risks of surveillance and to support efforts in resistance and refusal to its use.
@@ 2.1. Workplace Surveillance Technologies
Technological advancements since the late twentieth century have led to the age of digital surveillance, making it easier for employers to constantly monitor their workers both inside and outside the workplace (Introna, 2000; Vitak and Zimmer, 2023a; Ajunwa et al., 2017). WSTs have the ability to track and even infer a worker’s productivity, behavior, and personal information through digital activity monitoring, camera and audio monitoring, and localization and biophysical monitoring (Ajunwa et al., 2017; Ball, 2010; Cousineau et al., 2023). As AI and sensing technologies become more affordable and accessible, workplaces have introduced more intrusive forms of surveillance that often involve an assemblage of networked devices to generate continuous fine-grained data about their workers (Marx, 2002). The responsibility of workplace monitoring has also become more distributed in nature, where employees are expected to aid in their own surveillance through participation in workplace wellness programs and the use of productivity apps marketed towards improving their work and wellbeing (Ajunwa et al., 2017; Ball, 2010; Chowdhary et al., 2023; Kawakami et al., 2023; Adler et al., 2022; Saha and Iqbal, 2023; Ajunwa et al., 2016). The increased gigification of work, coupled with the rise of remote work during the COVID-19 pandemic, has seen a proliferation of these technologies being used beyond the workplace, further blurring the lines between work and private life (Calacci, 2022; Allyn, 2020; Satariano, 2020; Zickuhr, 2021; Gould et al., 2023).
@@
Although these proposed benefits highlight the potential for monitoring technologies to support workers, there have been numerous concerns that undermine this vision, including threats to worker privacy, autonomy, health, and safety, and the reinforcement of existing biases (Awumey et al., 2024; Constantinides and Quercia, 2022). These technologies often collect extraneous and irrelevant information that do not meaningfully correlate to attributes of success on the job (Roemmich et al., 2023; Pritchard et al., 2015). In particular, AI systems that capture and infer worker emotion, mood, and affect have been susceptible to inaccuracies and bias (Roemmich et al., 2023; Boyd and Andalibi, 2023; Corvite et al., 2023; Das Swain et al., 2023; Kaur et al., 2022). Workers have reported experiencing notable shifts in behavior and interpersonal dynamics, including increased stress, disciplinary action, distrust, and competition (Pritchard et al., 2015; Levy, 2015; Alimahomed-Wilson and Reese, 2021). Pritchard et al. (Pritchard et al., 2015) found that London bus drivers performed a form of “self-surveillance” that involved proactively altering their behavior to avoid being flagged by monitoring systems or to beat their previous performance scores. Power differentials and a lack of transparency make it difficult for workers to meaningfully consent to these technologies, as they have little power to opt out or be in a position to understand and control the extent of monitoring, their data, or how they are being assessed (Chowdhary et al., 2023; Sannon et al., 2022; Ajunwa, 2020; Ajunwa et al., 2016). In some cases, monitoring systems have led to unsafe working conditions and poorer performance, due to workers being compelled to perform additional labor in the form of continuously checking, negotiating, and responding to these systems (Pritchard et al., 2015; Lu et al., 2021; Sannon et al., 2022). Women and minority workers, in particular, have been disproportionately impacted as they are more likely to be employed in low-wage jobs and face biases that make them more susceptible to invasive monitoring (Ball, 2010; Stark et al., 2020; Anjali Anwar et al., 2021; Negrón and Nguyen, 2023).
@@
Workers have limited legal protections in the face of workplace surveillance (Zickuhr, 2021). In the US, there are currently no federal laws in place to protect against invasive workplace monitoring and legislation is limited in piecemeal form at the local and state levels (Ajunwa et al., 2017). The California Privacy Rights Act of 2020 is one of the few examples that offers some privacy protections to workers (Gray, 2020). Although frequently used as an exemplar for omnibus legislation that provides robust data protections, there remain gaps within the European Union’s General Data Protection Regulation (GDPR) that enables organizations to surveil workers while still being in compliance (Jodka, 2018; Horan, 2018). As a result, there have been calls for stronger organizational and governmental regulation to protect worker autonomy, privacy, and collective rights (Das Swain et al., 2023; Ebert et al., 2021; Nguyen, 2021; Calacci and Stein, 2023) and for technologists to take a worker-centered design approach that meaningfully engages with workers as collaborators, limits unnecessary data capture, and recognizes the nuanced and contextual aspects of their labor that are harder to measure or make visible (Calacci, 2022; Holten Møller et al., 2021; Zhang et al., 2022; Kawakami et al., 2023; Cameron et al., 2021; Akridge et al., 2024). Our work expands upon this literature by taking a holistic view on how workers experience and make sense of surveillance technologies across sectors, particularly amid the rise of remote work during the COVID-19 pandemic and shortly thereafter. We do so by drawing from worker testimonials on Reddit and in-depth interviews.
@@ 2.2. Worker Power and Resistance
Since the beginning of workplace surveillance, there has been worker resistance (Sale, 1995; Luff, 2008). Earlier studies of WSTs have shown how call center workers manipulate their metrics, such as leaving lines open after hanging up or pretending to talk on the phone (Frenkel et al., 1998; Ball, 2010; Callaghan and Thompson, 2002). Despite technologies becoming more sophisticated, workers have still found creative ways to circumvent them. For example, gig workers use sousveillance (Mann and Ferenbok, 2013; Mann et al., 2002; Cecchinato et al., 2021) techniques to monitor “up” by tracking payment discrepancies in ridehail work or collecting information about specific requesters on Amazon Mechanical Turk (Griffith, 2022; Hara et al., 2018; Do et al., 2024; Sannon et al., 2022). Additionally, workers use self-surveillance techniques by tracking their own personal data as a way to hold their employers and customers accountable (Sannon et al., 2022; Gallagher et al., 2023). “Worker inquiry” through data collection and analysis has long been an advocacy tool used by workers and unions since the days of Taylorism (Sannon et al., 2022; Khovanskaya et al., 2019; United Electrical and of America, 1943; Khovanskaya and Sengers, 2019; Dalal, 2024; Alquati, 2019). Sannon et al. (Sannon et al., 2022) broadly call these tactics self-protective surveillance that help to shield workers from risky work situations. They outline other resistance tactics gig workers used to protect their work and privacy, including using platforms like Reddit to commiserate with other workers, creating false personas, or using technologies like VPNs to protect their personal information (Sannon et al., 2022). Surveillance studies scholar Gary Marx outlines 11 resistance tactics in total that workers use to subvert surveillance, including discovery, avoidance, piggybacking, switching, distorting, blocking, masking, breaking, refusal, explaining and contesting, and cooperation (Marx, 2003). Other workers have collectively organized to form unions or participate in strikes to fight against overt surveillance and poor working conditions (Alimahomed-Wilson and Reese, 2021).
@@
To address these power asymmetries, some HCI researchers have investigated how technological tools can better support workers in these countermeasures (Hara et al., 2018; Salehi et al., 2015; Irani and Silberman, 2013; Calacci and Pentland, 2022; Zhang et al., 2023; Toxtli and Savage, 2023; Toxtli et al., 2021; Platform, [n. d.]; Cooperative, [n. d.]; Job, [n. d.]; Lee et al., 2015). One of the most well-known examples is Turkopticon, an online platform where Amazon Mechanical Turk workers can rate, review, and vet requestors. Calacci & Pentland (Calacci and Pentland, 2022) co-built the Shipt calculator with worker groups to allow gig workers to track and share wage data and surface pay disparities. While these two tools have helped support workers in collective data stewardship, sensemaking, and organizing, Calacci states that more can be done to support “Digital Workerism” by co-building tools with workers that support their goals in collective organizing while preserving their autonomy, privacy, and wellbeing (Calacci, 2022, 2023). While past HCI scholarship has mainly focused on the resistance tactics employed by gig workers, our paper examines the countermeasures used by workers across multiple types of employment, including operations, customer service, marketing, and food and beverage service.
@@ 3.1.1. Reddit as a Site of Study
In line with other studies that focus on sensitive topics such as mental health, addiction, and pregnancy loss, we chose to study forums on Reddit due to its support for anonymous disclosure (Ammari et al., 2018; Andalibi et al., 2016; Andalibi and Forte, 2018; van der Nagel and Frith, 2015; Gauthier et al., 2022), which allows people to discuss about sensitive or stigmatized topics without having to disclose their real identity (Newman et al., 2011; van der Nagel and Frith, 2015). Researchers also have looked to Reddit to understand the experiences of gig workers, since online communities have been crucial for gig workers to share knowledge and build community despite the distributed nature of their work (Sannon et al., 2022; Carlos Alvarez de la Vega et al., 2021; Ramesh et al., 2023). In particular, Sannon et al. (Sannon et al., 2022) used Reddit to understand gig workers’ experiences regarding privacy and surveillance. Expanding upon this work, we chose Reddit as a site to understand workers’ experiences with surveillance technologies across different sectors and how they used the platform to share knowledge and support one another.
@@ 3.1.2. Data Collection
We first identified posts that discussed workplace surveillance through manual keyword searches using surveillance-related terms such as “track,” “monitor,” and “surveil,” paired with keywords like “work” and “job.” We then randomly selected 66 posts to do a preliminary round of coding to find other relevant keywords, initial codes, and identify worker-centered subreddits to include in our study. We excluded subreddits that did not primarily focus on workers, (e.g. r/Privacy, r/Technology, r/legaladvice), subreddits that focus on managers or professional roles that do surveillance-type work (e.g. r/sysadmin, r/askmanagers), as well as industry or company-specific subreddits (e.g r/doordash_drivers, r/mturk, r/Nanny). This allowed us to understand how information was being shared cross-sectorally between workers in different industries and to get a sense of the breadth of different types of bossware and their commonalities and differences. Finally, we excluded 1 worker subreddit that had explicit policies against data collection. In total, we selected 9 subreddits, including 2 general work subreddits (r/jobs, r/work), 5 specific to remote work (r/overemployed, r/RemoteJobs, r/remotework, r/workfromhome, r/WorkOnline) and 2 that had a political lean (r/antiwork, r/WorkersStrikeBack). Based on our initial coding, we finalized our keyword list to the following: “track,” “monitor,” “surveil,” “spy,” “productivity,” “micromanage,” and “bossware.”
Since we were interested in understanding the impacts of COVID-19 on workplace surveillance, we only included posts written between 1/1/2019 and 12/31/2022. We also excluded posts that had less than average engagement, which was calculated by identifying the median number of comments per post according to each subreddit. Based on this criteria, we created our filtered dataset through Pushshift API (stuck_in_the_matrix et al., [n. d.]) , which contained a total of 9,362 filtered posts. For each post, we extracted the keywords used, their comments, date, title, username, URL, and text.